suppose two sides of a rectangle measure 28 cm and 840 mm what is the area of the rectangle
Answer by mathsmiles(68)

Aha! Tricky! They put the length and width in different metrics. First, we need to convert one of them to get the other.

28cm and 840mm
Let's convert cm to mm (because there will be more mm than cm we won't have to deal with decimal points).
28cm = 10*28mm = 280mm

Area of a rectangle formula is
A = L x W

A = 280 x 840
A = 235200 mm^2 (millimeter squared)
Now this looks weird. We should convert it back to cm
A = 235200/10 cm^2 = 23520 cm^2 But even that looks weirdly big:
A = 23520/10 dm^2 = 2352 dm^2 (decimeter squared)
Or go all out and convert to meters squared:
A = 235.2 m^2 (meters squared)